40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Feb 15 23:19:54 2011, Frank, DailyProgress, BEAT, some (minor) calibration problem found frequency_drift.pngfull_error_signal.png
    Reply  Tue Feb 15 23:33:51 2011, tara, DailyProgress, BEAT, some (minor) calibration problem found 
       Reply  Wed Feb 16 09:14:14 2011, Frank, DailyProgress, BEAT, some (minor) calibration problem found 
Message ID: 501     Entry time: Wed Feb 16 09:14:14 2011     In reply to: 499
Author: Frank 
Type: DailyProgress 
Category: BEAT 
Subject: some (minor) calibration problem found 

yes, for yesterdays measurements 7.435 MHz/V.

But that number changes from measurement to measurement as i'm changing the setup every time to improve SNR (or bring back stuff i've stolen from 40m)


So, from the first plot, is the calibration factor for cable delay technique ~ 8MHz/Volt?


Found at least two (minor) problems:

  1. So far i measured the zero-crossing of the mixer signal and both peak values and modeled the signal using a sine.
    For calibration i used the slope of the sine near the zero-crossing.
    As the frequency range for pi in shift is about 60MHz the function is almost linear for the range we measure over 24h (~2MHz).
    Today i measured the mixer signal in small steps and compared both techniques. The result surprised me a little bit:

    it is not a lot of difference in slope but it is surprisingly linear over a range of 40MHz!

  3. the channel which contains the calibrated VCO monitor signal uses a wrong sign.
    The fit is right, but i remembered that i exchanged the BNC to 2-pole LEMO connector some time ago because it had the wrong sign.
    However, the frequency tuning curve i measured with the old (wrong) cable.
    I used the right sign on my computer but implemented the wrong one in EPICS, so that data was/is wrong. Will change that tomorrow.
    This can't explain the problems i had last week as i used Matlab to convert the monitor signal into frequency, not that channel.
    Anyway, this is how the data i took last night (19h) looks like using the right slope (see item 1):
    so i think once the EPICS channel is fixed we can trust that channel



ELOG V3.1.3-